# AirRAG: Activating Intrinsic Reasoning for Retrieval Augmented Generation using Tree-based Search

### v2

@@ Abstract
Leveraging the autonomous decision-making capabilities of large language models (LLMs) has demonstrated superior performance in reasoning tasks. However, despite the success of iterative or recursive retrieval-augmented generation (RAG) techniques, these methods are often constrained to a single solution space when confronted with complex problems. In this paper, we propose a novel thinking pattern in RAG that integrates system analysis with efficient reasoning actions, significantly Activating intrinsic reasoning capabilities and expanding the solution space of specific tasks via Monte Carlo Tree Search (MCTS), which we refer to as AirRAG. Specifically, our approach designs five fundamental reasoning actions, which are expanded to a broad tree-based reasoning space using MCTS. The approach also incorporates self-consistency verification to explore potential reasoning paths and inference scaling law. Additionally, computationally optimal strategies are employed to allocate more inference resources to key actions, thereby enhancing overall performance. Experimental results demonstrate the effectiveness of AirRAG, showing significant performance gains on complex question-answering datasets. Furthermore, AirRAG is flexible and lightweight, making it easy to integrate with other advanced technologies.
AirRAG: Activating Intrinsic Reasoning for Retrieval Augmented Generation using Tree-based Search
Wenfeng Fengthanks: The first two authors contributed equally, Chuzhan Hao11footnotemark: 1, Yuewei Zhang, Jingyi Song & Hao Wangthanks: Corresponding author Alibaba Cloud Computing {wenfeng.fwf,haochuzhan.hcz,liyou.zyw,sjy267290}@alibaba-inc.com cashenry@126.com
@@ 1 Introduction
Previous research on complex query scenarios has primarily focused on optimizing the query and retrieval processes to obtain relevant information (Shi et al., 2023; Zhou et al., 2023; Gao et al., 2023; Jiang et al., 2023; Zheng et al., 2024). While these methods improve task efficiency, their performance often depends heavily on manually crafted rules and prompt engineering, limiting flexibility and making it difficult to adapt quickly to different scenarios. Moreover, recursive retrieval is frequently used to improve the depth and relevance of search results in information retrieval tasks. This process continuously updates intermediate queries and results to satisfy dynamic information needs during the complex task-solving process (Jiang et al., 2023; Asai et al., 2024; Yan et al., 2024; Jeong et al., 2024). Recent approaches leverage the decision-making abilities of LLMs to implement automated iterative retrieval and generation (Yu et al., 2024; Yue et al., 2024).
@@
However, these approaches face two significant issues. First, modular RAG struggles to cover the diverse range of question types, requiring substantial human effort and resulting in inefficiencies. The limited solution space also prevents the full activation of LLMs’ decision-making abilities. Second, the single reasoning paradigm and the chain-like reasoning process often fail to effectively explore the solution space, particularly when reasoning relies on self-exploration. This process is vulnerable to low-quality intermediate reasoning steps and is easily trapped in a narrow solution space. Furthermore, guiding self-exploration becomes challenging when using relatively smaller language models (e.g., Qwen2.5-14B-Instruct (Yang et al., 2024)). Therefore, it is necessary to design rational reasoning actions that comprehensively explore a broad and deep solution space.
@@
In response to these challenges, we propose AirRAG, a method that leverages intrinsic reasoning capabilities and expands the solution space using Monte Carlo Tree Search (MCTS). We design five fundamental reasoning actions: system analysis, direct answer, retrieval-answer, query transformation, and summary-answer. These actions can effectively address a wide range of problems in various scenarios, including those that require progressive or parallel queries. Importantly, these actions can be executed efficiently even on relatively smaller language models. Additionally, we introduce MCTS and self-consistency to enable controllable reasoning path generation and efficient inference scaling. To accurately select the answer from multiple reasoning paths, we combine a voting mechanism with a process-supervised reward model. As inference computation increases, our approach demonstrates significant performance improvements as shown in Figure 1. Moreover, AirRAG features a flexible architecture that can easily integrate other advanced methods into the approach as additional action branches. In summary, our main contributions are as follows:
@@
  • We design five fundamental reasoning actions that can address most problem types in RAG scenarios, ensuring controllable reasoning processes.
@@
  • We show through experimental results that AirRAG outperforms current iterative or recursive RAG approaches, effectively activating the intrinsic reasoning capabilities of LLMs and expanding the solution space in a controllable manner.
@@ 2 Related Work
Activating the reasoning capabilities of LLMs in RAG. In addition to optimizing retrieval, activating the reasoning capabilities of LLMs can significantly improve the efficiency and relevance of the retrieved information. Leveraging the decision-making abilities of LLMs enhances the overall performance (Nakano et al., 2022; Schick et al., 2023). Self-RAG and its variants (Asai et al., 2024; Yan et al., 2024; Jeong et al., 2024) adopt a self-reflection mechanism that iteratively predicts reflection tokens during training, enabling better control during inference. Auto-RAG (Yu et al., 2024) systematically plans retrievals and refines queries to acquire valuable knowledge through multi-turn iterations. IterDRAG (Yue et al., 2024) explores inference scaling strategies in RAG, enhancing LLMs’ ability to effectively acquire and utilize contextual information. Despite the progress made in these methods, they often struggle to explore the solution space effectively during reasoning. Self-exploration frequently leads to being trapped in a limited solution space, hindered by low-quality reasoning steps even after multiple iterations. This issue is often attributed to the chain reasoning pattern and the difficulty small-scale LLMs face when handling overly complex tasks in a single iteration.
@@
Unlike existing methods that focus on optimizing query and retrieval processes or leveraging LLMs’ reasoning capabilities through iterative retrieval, AirRAG uniquely integrates MCTS and self-consistency to systematically expand the solution space and ensure the controllability of the reasoning process. Simultaneously, we design five fundamental reasoning actions that effectively address a broader range of question types, particularly in complex tasks. This pluggable architecture also allows for easy integration of current advanced methods, making AirRAG a flexible and powerful solution for RAG scenarios. In the experiment, we thoroughly verify the performance gains brought by the inference scaling law and investigate how to rationally allocate inference resources.
@@ 3.1 Define Fundamental Reasoning Actions
Relying solely on the autonomy of LLMs for iterative self-exploration often results in getting trapped in a solution space that is difficult to navigate, especially when dealing with different types of complex questions. IterDRAG (Yue et al., 2024) uses a single action type to generate the next reasoning step, which can lead to ineffective space exploration. The core of MCTS generation lies in the action space, which defines the scope of tree exploration. Therefore, simplifying human cognitive processes in complex reasoning is essential (Jaffe et al., 2023). Inspired by this, we introduce five fundamental human-like reasoning actions to bridge the gap between LLM reasoning and human cognition in RAG scenarios.
@@
The above five actions define a highly diverse action space $\{A_{1},A_{2},A_{3},A_{4},A_{5}\}$. In the first step, the initial state is denoted as $s_{0}$ and then MCTS selects the action $a_{1}$ and $a_{2}$ to prompt the LLM to generate the next reasoning steps in parallel. Subsequent actions are performed sequentially to expand the reasoning path. It is important to note that there are sequential dependencies between different actions. For example, $A_{1}$ and $A_{2}$ can only be executed after the root question. Additionally, we incorporate the diverse sampling of self-consistency (Wang et al., 2023) for each action to expand the reasoning paths. Specifically, an action is more likely to generate the correct reasoning step if we sample multiple times in the current state. Finally, we can obtain multiple generated reasoning trajectories, such as $[s_{0}\oplus s_{1:n}]$. To further improve inference efficiency, we choose the action $\{A_{3},A_{4},A_{5}\}$ as a simplified action space, referred to as AirRAG-Lite, which achieves a better balance between efficiency and effectiveness.
@@ 3.2.1 Solution Generation
where $Q(s,a)$ is the reward value for node $s$ and is updated through backpropagation. $N(s)$ denotes the number of visits to $s$, $p$ is the parent node of $s$, and $w$ is the weight to balance exploration and exploitation.
@@ 3.2.3 Flexible Architecture
Our tree-based architecture provides the flexibility to integrate other advanced approaches. We reproduce the IterDRAG method based on the prompt design by Yue et al. (2024). Meanwhile, inspired by its iterative implementation, we simplify the fundamental action space to $\{A_{3},A_{4},A_{5}\}$, enabling a faster implementation while still achieving relatively good results. These methods serve as an exploratory extension of our approach and can be activated or deactivated as needed.
@@ 4 Experiments
  • RQ5: What is the intuitive performance of AirRAG in the reasoning process?

### v3

# AirRAG: Autonomous Strategic Planning and Reasoning Steer Retrieval Augmented Generation

@@ Abstract
Leveraging the autonomous decision-making capabilities of large language models (LLMs) has demonstrated superior performance in reasoning tasks. However, despite the success of iterative or agentic retrieval-augmented generation (RAG) techniques, these methods are often constrained to a single solution space when confronted with complex problems. In this paper, we propose a novel thinking pattern in RAG that integrates autonomous strategic planning with efficient reasoning actions, significantly activating intrinsic reasoning capabilities and expanding the solution space of specific tasks via Monte Carlo Tree Search (MCTS), which we refer to as AirRAG. Specifically, our approach designs five fundamental reasoning actions, which are expanded to a broad tree-based reasoning space using MCTS. The approach also incorporates self-consistency verification to explore potential reasoning paths and inference scaling law. Additionally, computationally optimal strategies are employed to allocate more inference resources to key actions, thereby enhancing overall performance. Experimental results demonstrate the effectiveness of AirRAG, showing significant performance gains on complex question-answering datasets. Furthermore, AirRAG is flexible and lightweight, making it easy to integrate with other advanced technologies and models.
AirRAG: Autonomous Strategic Planning and Reasoning Steer Retrieval Augmented Generation
Wenfeng Fengthanks: The first two authors contributed equally, Chuzhan Hao11footnotemark: 1, Yuewei Zhang, Guochao Jiang, Jingyi Song & Hao Wangthanks: Corresponding author Alibaba Cloud Computing {wenfeng.fwf,haochuzhan.hcz}@alibaba-inc.com,cashenry@126.com
@@ 1 Introduction
Previous research on complex query scenarios has primarily focused on optimizing the query and retrieval processes to obtain relevant information (Shi et al., 2023; Zhou et al., 2023; Gao et al., 2023; Jiang et al., 2023; Zheng et al., 2024; Asai et al., 2024; Yan et al., 2024). Iterative retrieval is frequently used to improve the depth and relevance of search results in information retrieval tasks. This process continuously updates intermediate queries and results to satisfy dynamic information needs during the complex task-solving process (Jeong et al., 2024; Yue et al., 2024). In addition, Li et al. (2025) integrates an agentic search workflow into the reasoning process, enabling dynamic retrieval when LLMs encounter uncertain knowledge points. The agentic LLMs are trained to learn step-by-step reasoning with search through reinforcement learning (Chen et al., 2025; Zheng et al., 2025).
@@
However, these approaches face two significant issues. First, the single reasoning paradigm and the chain-like reasoning process often fail to effectively explore the solution space, particularly when reasoning relies on self-exploration. This process is vulnerable to low-quality intermediate reasoning steps and is easily trapped in a narrow solution space. Second, the agentic search workflow and guiding self-exploration become challenging when using relatively smaller language models (e.g., Qwen2.5-7B-Instruct (Yang et al., 2024)). Furthermore, trainable agentic LLMs require efficient reinforcement learning training data and are difficult to apply to models with hundreds of billions of parameters.
@@
In response to these challenges, we propose AirRAG, a method that leverages autonomous strategic planning and reasoning capabilities and expands the solution space using Monte Carlo Tree Search (MCTS). We design five fundamental reasoning actions: system analysis, direct answer, retrieval-answer, query transformation, and summary-answer. These actions are the core and frequently used ones in the deep search scenarios, which can effectively address a wide range of problems in various scenarios, including those that require progressive or parallel queries. Importantly, these actions can be executed efficiently on LLMs of different scales. Additionally, we introduce MCTS and self-consistency to enable controllable reasoning path generation and efficient inference scaling. To accurately select the answer from multiple reasoning paths, we combine a voting mechanism with a process-supervised reward model. As inference computation increases, our approach demonstrates significant performance improvements as shown in Figure 1. Moreover, AirRAG features a flexible architecture that can easily integrate other advanced methods into the approach as additional action branches. In summary, our main contributions are as follows:
@@
  • We design five fundamental reasoning actions that can address most problem types in deep search scenarios, ensuring controllable reasoning processes.
@@
  • We show thorough experimental results that AirRAG outperforms current iterative or agentic methods, effectively activating the planning and reasoning capabilities of LLMs and flexibly expanding the solution space.
@@ 2 Related Work
Autonomous Planning and Reasoning in RAG. In addition to optimizing retrieval, activating the planning and reasoning capabilities of LLMs can significantly improve the efficiency and relevance of the retrieved information. Leveraging the decision-making abilities of LLMs enhances the overall performance (Nakano et al., 2022; Schick et al., 2023). Self-RAG and its variants (Asai et al., 2024; Yan et al., 2024; Jeong et al., 2024) adopt a self-reflection mechanism that iteratively predicts reflection tokens during training, enabling better control during inference. Auto-RAG (Yu et al., 2024) systematically plans retrievals and refines queries to acquire valuable knowledge through multi-turn iterations. IterDRAG (Yue et al., 2024) explores inference scaling strategies in RAG, improving LLMs’ ability to effectively acquire and utilize contextual information. Search-o1 (Li et al., 2025) designs an agentic search workflow to dynamically obtain effective knowledge. ReSearch (Chen et al., 2025) and DeepResearcher (Zheng et al., 2025) train agentic LLMs to reason with search using reinforcement learning. Despite the progress made in these methods, they often struggle to explore the solution space effectively during reasoning. Self-exploration frequently leads to being trapped in a limited solution space, hindered by low-quality reasoning steps even after multiple iterations. This issue is often attributed to the chain reasoning pattern and the difficulty small-scale LLMs face when handling overly complex tasks in a single iteration.
@@
Unlike existing methods that focus on optimizing query and retrieval processes or leveraging LLMs’ reasoning capabilities through iterative retrieval, AirRAG uniquely integrates MCTS and self-consistency to systematically expand the solution space and ensure the controllability of the reasoning process. Simultaneously, we design five fundamental reasoning actions that effectively address a broader range of question types, particularly in complex tasks. This pluggable architecture also allows for easy integration of current advanced methods or reasoning language models, making AirRAG a flexible and powerful solution for deep search scenarios. In the experiment, we thoroughly verify the performance gains brought by the inference scaling law and investigate how to rationally allocate inference resources.
@@ 3.1 Define Fundamental Reasoning Actions
Relying solely on the autonomy of LLMs for iterative self-exploration often results in getting trapped in a solution space that is difficult to navigate, especially when dealing with different types of complex questions. IterDRAG (Yue et al., 2024) uses a single action type to generate the next reasoning step, which can lead to ineffective space exploration. The core of MCTS generation lies in the action space, which defines the scope of tree exploration. Based on advanced methods and reasoning language models, we summarize the most common actions in RAG, such as query transformation and retrieval answering. Meanwhile, the chain-of-thought in reasoning models has demonstrated superior performance in complex open-domain question answering. Therefore, simplifying human cognitive processes in complex reasoning is essential (Jaffe et al., 2023). Inspired by this, we introduce five fundamental human-like reasoning actions to bridge the gap between LLM reasoning and human cognition in RAG scenarios.
@@
The above five actions define a highly diverse action space $\{A_{1},A_{2},A_{3},A_{4},A_{5}\}$. In the first step, the initial state is denoted as $s_{0}$ and then MCTS selects the action $a_{1}$ and $a_{2}$ to prompt the LLM to generate the next reasoning steps in parallel. Subsequent actions are performed sequentially to expand the reasoning path. It is important to note that there are sequential dependencies between different actions. For example, $A_{1}$ and $A_{2}$ can only be executed after the root question. Additionally, we incorporate the diverse sampling of self-consistency (Wang et al., 2023) for each action to expand the reasoning paths. Specifically, an action is more likely to generate the correct reasoning step if we sample multiple times in the current state. Finally, we can obtain multiple generated reasoning trajectories, such as $s_{0}\oplus s_{1:n}]$. To further improve inference efficiency, we choose the action $\{A_{3},A_{4},A_{5}\}$ as a simplified action space, referred to as AirRAG-Lite, which achieves a better balance between efficiency and effectiveness.
@@ 3.2.1 Solution Generation
where $Q(s,a)$ is the reward value for node $s$, generated by taking action $a$, and is updated through backpropagation. $Q(s,a)$ serves as a key metric in MCTS to evaluate the value of each node. $N(s)$ denotes the number of visits to $s$, $p$ is the parent node of $s$, and $w$ is the weight to balance exploration and exploitation. Initially, all unexplored nodes are assigned $Q(s_{i},a_{i})=0$, leading to random tree expansions at the beginning. When the search reaches a terminal node $n_{d}$, a reward score $Q(s_{d},a_{d})$ is computed based on whether the node reaches the correct answer. This score is then back-propagated to all intermediate nodes along the trajectory $t=x\oplus s_{1}\oplus s_{2}\oplus\ldots\oplus s_{d}$. Specifically, for each intermediate node $s_{i}$ (for $i=1,2,\ldots,d-1$), its $Q$ value is updated as: $Q(s_{i},a_{i})=Q(s_{i},a_{i})+Q(s_{d},a_{d})$.
@@ 3.2.3 Flexible Architecture
Our tree-based architecture provides the flexibility to integrate other advanced approaches. We reproduce the IterDRAG method based on the prompt design by Yue et al. (2024). Meanwhile, inspired by its iterative implementation, we simplify the fundamental action space to $\{A_{3},A_{4},A_{5}\}$, enabling a faster implementation while still achieving relatively good results. These methods serve as an exploratory extension of our approach and can be activated or deactivated as needed. Due to the training-free nature of our method, its generator LLM can be arbitrarily replaced with the strongest existing models for performance improvement.
@@ 4 Experiments
  • RQ5: What is the intuitive efficiency and performance of AirRAG in the reasoning process?